OpenStudy (anonymous):

Multiply three-eighths times twelve. one thirty-second 4 and one-half thirty-two 1 and seven-eighths

OpenStudy (anonymous):

3/8 x 12 = 3/2 x 3 (Divide both 8 and 12 by 4) = 9/2 = 4 and one half I hope this helps.